OpenStudy (bbb911):

This is Geometry: Solve by using substitution or elimination: x=3y+11 2x+5y=0

OpenStudy (anonymous):

Definite substitution. PLug the first equation into the 2nd in place of x. Solve for y then plug back into the 1st to get x.

OpenStudy (anonymous):

2(3y + 11) + 5y = 0 6y + 22 + 5y = 0 11y + 22 = 0 11y = -22 y = -2 Now get x...

OpenStudy (bbb911):

thank you for helping :3

OpenStudy (anonymous):

my pleasure